Assume that the banking system has total reserves of Rs.150 billion. Assume also that required reserves are 20 percent of checking deposits and that banks hold no excess reserves and households hold no currency.

  1. Calculate the money multiplier? 
  2. Calculate the money supply? 

If the State Bank of Pakistan now raises required reserves to 25 percent of deposits,

  1. Calculate the money multiplier? 
  2. What will be the effect on Reserves?  (Please write only one word “Increase”, “Decrease”, or “No Change” in the blank)
  3. The amount of money supply will decline to 
1
Expert's answer
2021-06-03T13:49:52-0400

(a)

1.

"money\space multiplier=\frac{1}{Reserve \space ratio}"

reserve ratio = % deposits


"\therefore money \space multiplier=\frac{1}{0.2}=5"



2.

money supply "=5\times 150\times 0.8=600"

therefore money supply is Rs. 600 billion


(b)

when required reserves are raised to 25 percent of deposits

1.


"money\space multiplier=\frac{1}{Reserve \space ratio}"


reserve ratio = % deposits


"\therefore money \space multiplier=\frac{1}{0.25}=4"



2. Increase


3.

money supply "=4\times 150\times 0.75=450"

therefore money supply is Rs. 450 billion